OpenStudy (anonymous):

A race car travels at a speed of 230 miles per hour. The race track on which the car is competing is 400 miles long. About how long does it take for the car to finish, disregarding pit stops and yellow flags? Step by step explanation would be nice, if possible. Thank you !

OpenStudy (anonymous):

were gonna use the formula for speed which is S=D/T... S is speed, D is distance, T is time. S=230miles per hour D=400miles T=? 230=400/T T=400/230 T=1.74hours

OpenStudy (anonymous):

Thank you,I really appreciate the help :]

OpenStudy (anonymous):

no probz.. =^_^=